OpenStudy (ccastudentgrade10):

WILL FAN AND MEDAL write the expression as a single logarithm 5logby + 4logbx i have no idea how to do this answers posted in a sec

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

no need to post the answers

OpenStudy (misty1212):

you will use two facts

OpenStudy (ccastudentgrade10):

hi! :)

OpenStudy (misty1212):

first one is that \[n\log(x)=\log(x^n)\] write each like that first

OpenStudy (misty1212):

second is that \[\log(x)+\log(y)=\log(xy)\] use that second

OpenStudy (ccastudentgrade10):

blog(y) = log(x^4) ?

OpenStudy (misty1212):

no don't set them equal first rewrite \[5\log(y)\] as \[\log(y^5)\] now you do the second one \[4\log(x)\]

OpenStudy (ccastudentgrade10):

log(x^4)

OpenStudy (misty1212):

right so now we are at this step \[5\log(y)+4\log(x)=\log(y^5)+\log(x^4)\]

OpenStudy (misty1212):

now the sum of the logs is the log of the products, which is an english way of saying \[\log(A)+\log(B)=\log(AB)\]

OpenStudy (misty1212):

so write the log of the product of \(y^5\) and \(x^4\)

OpenStudy (ccastudentgrade10):

log(xy^9) ?

OpenStudy (ccastudentgrade10):

@misty1212

OpenStudy (misty1212):

no

OpenStudy (misty1212):

\[y^5x^4\neq xy^9\]

OpenStudy (misty1212):

\[x^4y^5=x^4y^5\]that is all

OpenStudy (misty1212):

so put that inside the log

OpenStudy (ccastudentgrade10):

i dont understand how

OpenStudy (ccastudentgrade10):

@misty1212

OpenStudy (misty1212):

just write it \[\log(x^4y^5)\]

OpenStudy (ccastudentgrade10):

so is that the answer? @misty1212

OpenStudy (anonymous):

yes

OpenStudy (ccastudentgrade10):

so it would be b, right?

OpenStudy (anonymous):

post your options

OpenStudy (anonymous):

yes

OpenStudy (ccastudentgrade10):

thank you!!

OpenStudy (anonymous):

np